### v1

# Left Jacobson rings

Abstract.
We say that a ring is strongly (resp. weakly) left Jacobson if every semiprime (resp. prime) left ideal is an intersection of maximal left ideals. There exist Jacobson rings that are not weakly left Jacobson, e.g. the Weyl algebra. Our main result is the following one-sided noncommutative Nullstellensatz: For any finite-dimensional $\mathbbm{F}$-algebra $\mathbbm{A}$ the ring $\mathbbm{A}[x_{1},\dots,x_{n}]$ of polynomials with coefficients in $\mathbbm{A}$ is strongly left Jacobson and every maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ has a special form. We also prove that an algebra that is a finitely generated module over its center is weakly left Jacobson iff it is Jacobson, and that an Azumaya algebra is strongly left Jacobson iff its center is Jacobson.
The first author is affiliated with the Faculty of Math. and Phys., University of Ljubljana, Slovenia, and with IMFM (Institute of Math., Phys. and Mech.)
The second author is affiliated with the Faculty of Math. and Phys., University of Ljubljana, Slovenia.
The first author acknowledges the support from grants P1-0222, J1-50002 and J1-60011 which are funded by ARIS (Slovenian Research and Innovation Agency)
[…]
1. Summary
A commutative ring is Jacobson if every prime ideal is an intersection of maximal ideals. By [12] or [9, Thm. 4.9], every finitely generated commutative $\mathbbm{F}$-algebra (where $\mathbbm{F}$ is a field) is Jacobson. This definition can be extended to non-commutative rings in different ways.
In [17, Chap. 9], an (associative unital) ring is called Jacobson if every prime ideal is equal to the intersection of all primitive ideals that contain it. Examples of Jacobson $\mathbbm{F}$-algebras are Weyl algebras $A_{n}(\mathbbm{F})$, enveloping algebras of finite-dimensional $\mathbbm{F}$-Lie algebras and group algebras $\mathbbm{F}[G]$ where $G$ is polycyclic-by-finite; see [17, Cor. 9.1.8]. Another example are finitely generated $\mathbbm{F}$-algebras that satisfy a polynomial identity; see [2, Cor. 1.3]. By [15, Thm. 10.11], every semiprime ideal is an intersection of prime ideals, so a ring is Jacobson if and only if every semiprime ideal equals the intersection of all primitive ideals that contain it.
We are interested in a different extension. We say that a ring is weakly (resp. strongly) left Jacobson if every prime (resp. semiprime) left ideal is equal to the intersection of all maximal left ideals that contain it. We use the definitions of prime and semiprime left ideals from [13, Sec. 1]. Note that every strongly left Jacobson ring is also weakly left Jacobson, and every weakly left Jacobson ring is Jacobson. Rings in which every semiprime left ideal is equal to the intersection of all prime left ideals that contain it will be called left intersection rings. A ring is strongly left Jacobson iff it is both weakly left Jacobson and a left intersection ring.
The aim of this paper is to provide new examples of strongly left Jacobson rings. Our main result is that any ring of polynomials with coefficients in a finite-dimensional $\mathbbm{F}$-algebra is strongly left Jacobson; see Theorem 27. Special cases of this result were known before, namely for the matrix algebras $\mathrm{M}_{n}(R)$ and the standard quaternion algebra over $R$, where $R$ is a commutative Jacobson ring; see [5, Thm. 3] and [1], [6, Thm. B]. We will also give examples of Jacobson rings that are not even weakly left Jacobson. The main examples are the Weyl algebra $A_{1}(\mathbbm{F})$ and the universal enveloping algebra of the Heisenberg-Lie algebra; see Section 3.
The paper consists of 6 sections. In Section 2 we provide geometric interpretations of Jacobson and weakly/strongly left Jacobson rings which differ in the definition of a “point”. In Section 3 we give examples of Jacobson rings that are not weakly left Jacobson. In Sections 4 and 5 we develop sufficient conditions for a ring to be weakly left Jacobson and strongly left Jacobson, respectively. In Section 6 we apply these results to polynomials with coefficients in finite-dimensional algebras and prove the main Theorem 27.
2. Geometric interpretation
We assume that all rings and algebras are associative and unital. Two-sided ideals will simply be called ideals.
For every ideal $I$ of a ring $R$ we define its Jacobson radical $\operatorname{rad}(I)$ as the intersection of all maximal left ideals that contain it. Equivalently, $\operatorname{rad}(I)$ is the intersection of all left-primitive ideals that contain $I$; see [15, Cor. 11.5]. By definition, an ideal is left-primitive iff it is equal to the annihilator of a simple left $R$-module. As usual we write $\operatorname{rad}R:=\operatorname{rad}(\{0\})$.
If $R$ is an algebra over a field $\mathbbm{F}$ then left $R$-modules correspond to representations of $R$, i.e. to $\mathbbm{F}$-algebra homomorphisms $\pi\colon R\to\mathrm{End}(V_{\pi})$ where $V_{\pi}$ is a vector space over $\mathbbm{F}$. Simple left modules correspond to irreducible representations and the annihilators of modules correspond to the kernels of representations. Therefore, left-primitive ideals coincide with the kernels of irreducible representations.
For every ideal $I$ of a ring $R$ we define its prime radical $\sqrt{I}$ as the intersection of all prime ideals that contain $I$. By [15, Cor. 10.12], $\sqrt{I}$ is the smallest semiprime ideal containing $I$. Since every primitive ideal is prime by [15, Cor. 11.6] we have $\sqrt{I}\subseteq\operatorname{rad}(I)$.
Note that $R$ is a Jacobson ring if and only if $\operatorname{rad}(I)=\sqrt{I}$ for every ideal $I$ of $R$. To give this observation a geometric interpretation in the spirit of Hilbert’s Nullstellensatz we introduce the following terminology. The elements of $R$ will be called “polynomials” and the simple left $R$-modules $M$ (or irreducible representations $\pi$ in the case of $\mathbbm{F}$-algebras) will be called “points”. We say that a “point” $M$ (resp. $\pi$) is annihilated by a “polynomial” $a$ if $aM=0$ (resp. $\pi(a)=0$). Proposition 1 summarizes the discussion.
Proposition 1.
A ring (or $\mathbbm{F}$-algebra) $R$ is Jacobson if and only if for every subset $G$ of $R$ and every “polynomial” $f\in R$ the following are equivalent:
  1. i)
    $f$ belongs to the smallest semiprime ideal that contains $G$.
  2. ii)
    Every “point” of $R$ which is annihilated by all $g\in G$ is also annihilated by $f$.
Remark 2.
The Jacobson property of an $\mathbbm{F}$-algebra $R$ can be considered a weak version of a noncommutative Hilbert Nullstellensatz for $R$. To obtain a proper Nullstellensatz we must reduce the set of “points” in claim ii) of Proposition 1 as much as possible. For commutative finitely generated $\mathbbm{F}$-algebras we can take (by Hilbert’s Nullstellensatz) $\mathbbm{F}$-algebra homomorphisms from $R$ into all finite field extensions of $\mathbbm{F}$, or equivalently into the algebraic closure of $\mathbbm{F}$; see [12, p. 140]. For the $\mathbbm{F}$-algebras mentioned in [17, Cor. 9.1.8] (i.e. Weyl algebras $A_{n}(\mathbbm{F})$, enveloping algebras of finite-dimensional $\mathbbm{F}$-Lie algebras and group algebras $\mathbbm{F}[G]$ where $G$ is polycyclic-by-finite) we can take all simple left $R$-modules for which $\mathrm{End}(M)$ is algebraic over $\mathbbm{F}$. For finitely generated $\mathbbm{F}$-algebras that satisfy a polynomial identity we can take finite-dimensional irreducible representations; see [2, Cor. 1.3].
Let us extend the definitions of Jacobson and prime radicals to one-sided ideals. We will use the definitions of prime and semiprime left ideals from [13, Sec. 1]. We say that a left ideal $I$ of a ring $R$ is prime if for every $a,b\in R$ such that $aRb\subseteq I$ we have $a\in I$ or $b\in I$. We say that $I$ is semiprime if for every $a\in R$ such that $aRa\subseteq I$ we have $a\in I$. Note that every prime left ideal is semiprime, and any intersection of semiprime left ideals is again semiprime.
[…]
Lemma 3 gives a geometric interpretation of maximal left ideals.
Lemma 3.
Let $R$ be a ring, $M$ a simple left $R$-module and $m$ a nonzero element of $M$. Then the annihilator $\operatorname{ann}_{M}(m)\coloneqq\{r\in R\mid rm=0\}$ of $m$ is a maximal left ideal of $R$. Moreover, every maximal left ideal of $R$ is of this form.
In particular, if $R$ is an $\mathbbm{F}$-algebra, then maximal left ideals of $R$ coincide with left ideals of the form $J_{\pi,v}\coloneqq\{a\in R\mid\pi(a)v=0\}$ where $\pi\colon R\to\mathrm{End}(V_{\pi})$ is an irreducible representation of $R$ and $v\in V_{\pi}\setminus\{0\}$.
Proof.
The annihilator $\operatorname{ann}_{M}(m)$ is clearly a left ideal of $R$. To show that it is maximal we have to prove that for every $c\in R\setminus\operatorname{ann}_{M}(m)$, $\operatorname{ann}_{M}(m)+Rc=R$. If $c\in R\setminus\operatorname{ann}_{M}(m)$, then $cm\neq 0$ which implies that $Rcm=M$ since $M$ is simple. In particular, there exist $r\in R$ such that $m=rcm$; i.e. $1-rc\in\operatorname{ann}_{M}(m)$.
To prove the converse pick a maximal left ideal $I$ of $R$ and note that $M\coloneqq R/I$ is a simple left $R$-module containing $m\coloneqq 1+I$. Clearly, $\operatorname{ann}_{M}(m)=I$. ∎
Let us define a “directional point” on $R$ as a pair $(M,m)$ where $M$ is a simple left $R$-module and $m\in M\setminus\{0\}$ (or in the case of $\mathbbm{F}$-algebras as a pair $(\pi,v)$ where $\pi\colon R\to\mathrm{End}(V_{\pi})$ is an irreducible representation of $R$ and $v\in V_{\pi}\setminus\{0\}$). We say that a “directional point” $(M,m)$ (resp. $(\pi,v)$) is annihilated by a “polynomial” $a\in R$ if $am=0$ (resp. $\pi(a)v=0$). Lemma 3 says that left maximal ideals of $R$ can be considered as the “annihilation ideals” of “directional points” on $R$.
In summary, we get the following geometric interpretation of the strong left Jacobson property (analogously for the weak version).
Proposition 4.
A ring (or $\mathbbm{F}$-algebra) $R$ is strongly left Jacobson if and only if for every subset $G$ of $R$ and every “polynomial” $f\in R$ the following are equivalent:
  1. i)
    $f$ belongs to the smallest semiprime left ideal that contains $G$.
  2. ii)
    Every “directional point” of $R$ which is annihilated by all $g\in G$ is also annihilated by $f$.
As in Remark 2 we will try for specific strongly left Jacobson $\mathbbm{F}$-algebras to reduce the set of “directional points” in part ii) of Proposition 4; e.g. to directional points $(\pi,v)$ where $\pi$ is finite-dimensional.
3. Counterexamples
The aim of this section is to give examples of $\mathbbm{F}$-algebras that satisfy the Jacobson property but do not satisfy the weak left Jacobson property.
Theorem 5.
For any field $\mathbbm{F}$ of characteristic 00 the first Weyl algebra
(1) $A_{1}(\mathbbm{F})=\mathbbm{F}\langle x,y\rangle/(yx-xy-1)$
is not weakly left Jacobson.
Proof.
We show first that every left ideal $I$ of $\mathbbm{A}$ is prime, so $I=\sqrt[wk]{I}$. Namely, if $a,b\in\mathbbm{A}$ fulfil $a\mathbbm{A}b\subseteq I$, then also $(\sum\mathbbm{A}a\mathbbm{A})b\subseteq I$. Since $\mathbbm{A}$ is simple we have either $\sum\mathbbm{A}a\mathbbm{A}=0$ or $\sum\mathbbm{A}a\mathbbm{A}=\mathbbm{A}$. In the first case $a=0\in I$, and in the second case $1\in\sum\mathbbm{A}a\mathbbm{A}$ implies $b\in I$.
We now take the left ideal $I\coloneqq\mathbbm{A}yx$. Clearly $x\notin I$. To show that $I\neq\operatorname{rad}(I)$ it suffices to show that $x\in\mathfrak{m}$ for every maximal left ideal $\mathfrak{m}$ of $\mathbbm{A}$ with $yx\in\mathfrak{m}$. So assume to the contrary that $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}$ and that $yx\in\mathfrak{m}$ but $x\notin\mathfrak{m}$. By maximality of $\mathfrak{m}$ there exists $r\in\mathbbm{A}$ such that $1+rx\in\mathfrak{m}$. We can expand $r$ as $r=\sum_{i=0}^{\ell}a_{i}(x)y^{i}\in\mathbbm{A}$ with suitable $\ell\in\mathbbm{N}_{0}$ and $a_{i}(x)\in\mathbbm{F}[x]$ for $i\in\{0,\dots,\ell\}$. Note that $yx^{k}-x^{k}y=kx^{k-1}$ for every $k\in\mathbbm{N}$, and $yx^{0}-x^{0}y=0$. Write $r^{(1)}\coloneqq yr-ry=\sum_{i=0}^{\ell}a_{i}^{(1)}(x)y^{i}$ where $a_{i}^{(1)}(x)$ is the first derivative of $a_{i}(x)$. Using analogous notation for arbitrary derivatives, we claim that $y^{n}+r^{(n)}x\in\mathfrak{m}$ for every $n\in\mathbbm{N}_{0}$. The case $n=0$ follows from the definition of $r$. Suppose now that the claim is true for some $n\in\mathbbm{N}_{0}$, then $y^{n+1}+r^{(n+1)}x=y^{n+1}+(yr^{(n)}-r^{(n)}y)x=y(y^{n}+r^{(n)}x)-r^{(n)}yx\in\mathfrak{m}$ since $\mathfrak{m}$ is a left ideal containing $yx$; so the claim is also true for $n+1$. If $n$ is larger than the maximum of degrees of $a_{i}(x)$, the claim implies that $y^{n}\in\mathfrak{m}$.
So we know that there exists $n\in\mathbbm{N}_{0}$ such that $y^{n}\in\mathfrak{m}$. We now show that in fact $y^{n}\in\mathfrak{m}$ for all $n\in\mathbbm{N}_{0}$. For this it remains to show that for any $k\in\mathbbm{N}$ such that $y^{k}\in\mathfrak{m}$ we also have $y^{k-1}\in\mathfrak{m}$. Namely, $ky^{k-1}=y^{k}x-xy^{k}=y^{k-1}yx-xy^{k}\in\mathfrak{m}$ since $yx\in\mathfrak{m}$ and $y^{k}\in\mathfrak{m}$; as $\mathbbm{F}$ has characteristic 00 this shows that indeed $y^{k-1}\in\mathfrak{m}$. In particular $1\in\mathfrak{m}$, so we have obtained a contradiction. ∎
To deduce other counterexamples we need the following:
Lemma 6.
Let $\pi\colon\mathbbm{A}\to\mathbbm{B}$ be a surjective ring homomorphism. The image and preimage under $\pi$ give a bijective correspondence between the left ideals of $\mathbbm{B}$ and the left ideals of $\mathbbm{A}$ that contain $\ker\pi$. It also preserves semiprime, prime and maximal left ideals. In particular if $\mathbbm{A}$ is weakly or strongly left Jacobson, then so is $\mathbbm{B}$.
Proof.
It is easy to check that the image and preimage under $\pi$ give a bijective correspondence between the left ideals of $\mathbbm{B}$ and the left ideals of $\mathbbm{A}$ that contain $\ker\pi$. This correspondence clearly preserves the order by inclusion, so it preserves maximal left ideals. In the following we check that this correspondence also preserves semiprime left ideals; the argument for prime left ideals is analogous.
Let $J$ be a semiprime left ideal of $\mathbbm{B}$. To prove that the left ideal $\pi^{-1}(J)$ is semiprime, pick $a\in\mathbbm{A}$ such that $a\mathbbm{A}a\subseteq\pi^{-1}(J)$. It follows that $\pi(a)\pi(\mathbbm{A})\pi(a)\subseteq J$. Since $J$ is semiprime and $\pi(\mathbbm{A})=\mathbbm{B}$, it follows that $\pi(a)\in J$. Therefore $a\in\pi^{-1}(J)$.
Let $I$ be a semiprime left ideal of $\mathbbm{A}$ containing $\ker\pi$. To show that $\pi(I)$ is semiprime, pick $b\in\mathbbm{B}$ such that $b\mathbbm{B}b\subseteq\pi(I)$. Since $\pi$ is surjective, we have $\mathbbm{B}=\pi(\mathbbm{A})$ and $b=\pi(a)$ for some $a\in\mathbbm{A}$. It follows that $a\mathbbm{A}a\subseteq\pi^{-1}(\pi(I))=I$ since $\ker\pi\subseteq I$, and so $a\in I$ since $I$ is semiprime. Therefore $b\in\pi(I)$.
Finally assume that $\mathbbm{A}$ is weakly (strongly) left Jacobson and let $J$ be a semiprime (prime) left ideal of $\mathbbm{B}$. Then $\pi^{-1}(J)$ is a semiprime (prime) left ideal of $\mathbbm{A}$, so $\pi^{-1}(J)=\bigcap M$ where $M$ is a set of maximal left ideals of $\mathbbm{A}$. For every $\mathfrak{m}\in M$ clearly $\ker\pi\subseteq\pi^{-1}(J)\subseteq\mathfrak{m}$, so $\pi(\mathfrak{m})$ is a maximal left ideal of $\mathbbm{B}$. It follows that $J=\bigcap\{\pi(\mathfrak{m})\mid\mathfrak{m}\in M\}$. This shows that $\mathbbm{B}$ is weakly (strongly) left Jacobson. ∎
As a special case of Lemma 6 we obtain:
Corollary 7.
Any algebra over a field of characteristic 00 which has $A_{1}(\mathbbm{F})$ as a homomorphic image is not weakly left Jacobson.
Let us give some special cases of Corollary 7.
Example 8.
The Heisenberg Lie algebra $\mathfrak{h}$ is the Lie algebra over a field $\mathbbm{F}$ with basis $X,Y,Z\in\mathfrak{h}$ and with the Lie bracket satisfying $[X,Y]=Z$ and $[X,Z]=[Y,Z]=0$. We have a homomorphism $\Phi$ from the enveloping algebra $U(\mathfrak{h})$ onto the Weyl algebra $A_{1}(\mathbbm{F})$ defined by $\Phi(X)=x$, $\Phi(Y)=y$ and $\Phi(Z)=-1$. If $\mathbbm{F}$ has characteristic 00, then $U(\mathfrak{h})$ is not left Jacobson.
Example 9.
Since $A_{1}(\mathbbm{F})=\mathbbm{F}\langle x,y\rangle/(yx-xy-1)$ is a homomorphic image of $\mathbbm{F}\langle x,y\rangle$, the $\mathbbm{F}$-algebra $\mathbbm{F}\langle x,y\rangle$ is not left Jacobson if $\mathbbm{F}$ has characteristic 00.
Recall that by [17, Cor 9.1.8] the algebras from Theorem 5 and Example 8 are Jacobson.
4. Some weakly Jacobson rings
In this section we show that a ring that is finitely generated as a module over its center is weakly left Jacobson if and only if it is Jacobson; see Corollary 15.
Consider any left ideal $I$ of a ring $\mathbbm{A}$ with center $R\coloneqq\mathcal{Z}(\mathbbm{A})$. Set
(2) $[I:\mathbbm{A}]\coloneqq\{a\in\mathbbm{A}\mid a\mathbbm{A}\subseteq I\}.$
This is a particular example of an ideal quotient. It is easy to check that $[I:\mathbbm{A}]$ is a two-sided ideal of $\mathbbm{A}$, and $[I:\mathbbm{A}]\subseteq I$ because $\mathbbm{A}$ has a unit.
Proposition 10.
Let $\mathbbm{A}$ be a ring with center $R\coloneqq\mathcal{Z}(\mathbbm{A})$ and $I$ a prime left ideal of $\mathbbm{A}$. Then $[I:\mathbbm{A}]$ is a (two-sided) prime ideal of $\mathbbm{A}$ and $I\cap R$ is a prime ideal of $R$. Moreover, $I\cap R=[I:\mathbbm{A}]\cap R$, which implies that the inclusion mapping $R\subseteq\mathbbm{A}$ induces an injective mapping of the integral domain $R/(I\cap R)$ into the center of the prime ring $\mathbbm{A}/[I:\mathbbm{A}]$.
Proof.
Consider $a,a^{\prime}\in\mathbbm{A}$ such that $a\mathbbm{A}a^{\prime}\subseteq[I:\mathbbm{A}]$. If $a^{\prime}\notin[I:\mathbbm{A}]$, then there exists $b^{\prime}\in\mathbbm{A}$ such that $a^{\prime}b^{\prime}\notin I$. However, $abca^{\prime}b^{\prime}\in I$ for all $b,c\in\mathbbm{A}$ because $abca^{\prime}\in[I:\mathbbm{A}]$, and therefore $ab\in I$ for all $b\in\mathbbm{A}$ because $I$ is prime, i.e. $a\in[I:\mathbbm{A}]$. This shows that $[I:\mathbbm{A}]$ is a prime ideal of $\mathbbm{A}$.
Next consider $r,s\in R$ such that $rs\in I$. Then $r\mathbbm{A}s=\mathbbm{A}rs\subseteq I$, so $r\in I$ or $s\in I$ because $I$ is prime. This shows that $I\cap R$ is a prime ideal of $R$.
Since $[I:\mathbbm{A}]\subseteq I$, we have $[I:\mathbbm{A}]\cap R\subseteq I\cap R$. To prove the converse pick $r\in I\cap R$ and note that $r\mathbbm{A}=\mathbbm{A}r\subseteq I$ which implies that $r\in[I:\mathbbm{A}]$. ∎
Definition 11.
Let $\mathbbm{A}$ be a ring with center $R\coloneqq\mathcal{Z}(\mathbbm{A})$ and $I$ a prime left ideal of $\mathbbm{A}$. By Proposition 10 we can identify the integral domain $\bar{R}\coloneqq R/(I\cap R)$ with a subset of the center of the prime ring $\bar{\mathbbm{A}}\coloneqq\mathbbm{A}/[I:\mathbbm{A}]$. Write $S\coloneqq\bar{R}\setminus\{0\}$ and note that the field $R_{I}\coloneqq S^{-1}\bar{R}$ is contained in the center of the prime ring $\mathbbm{A}_{I}\coloneqq S^{-1}\bar{\mathbbm{A}}$. For every $a\in\mathbbm{A}$ and $q\in R\setminus I$ we write $a/q\coloneqq(\bar{q})^{-1}\bar{a}\in\mathbbm{A}_{I}$ where $\bar{a}$ is the projection of $a$ to $\bar{\mathbbm{A}}$ and $\bar{q}$ is the projection of $q$ to $\bar{R}$.
Remark 12.
If $a,b\in\mathbbm{A}$ and $q,s\in R\setminus I$ then $a/q+b/s=(sa+qb)/(qs)$ and $(a/q)(b/s)=(ab)/(qs)$. Moreover, $a/q=b/s$ iff $sa-qb\in[I:\mathbbm{A}]$. To see that $\mathbbm{A}_{I}$ is indeed prime assume that $(a/q)(x/1)(b/s)=0/1$ for every $x\in\mathbbm{A}$. It follows that $axb\in[I:\mathbbm{A}]$ for all $x\in\mathbbm{A}$. Since $[I:\mathbbm{A}]$ is a prime ideal, $a\in[I:\mathbbm{A}]$ or $b\in[I:\mathbbm{A}]$. So $a/q=0/1$ or $b/s=0/1$.
Proposition 13.
Let $\mathbbm{A}$ be a ring with center $R\coloneqq\mathcal{Z}(\mathbbm{A})$ and $I$ a prime left ideal of $\mathbbm{A}$. Assume $I\neq\mathbbm{A}$. Then
(3) $\langle I\rangle\coloneqq{}$$\bigl{\{}a/q\mid a\in I\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ q\in R\setminus I\bigr{\}}$
is a left ideal of $\mathbbm{A}_{I}$ and
(4) $I={}$$\bigl{\{}a\in\mathbbm{A}\mid a/1\in\langle I\rangle\bigr{\}}.$
Proof.
Given $a/q,b/s\in\langle I\rangle$ with $a,b\in I$ and $q,s\in R\setminus I$, then $a/q+b/s=(sa+qb)/(qs)\in\langle I\rangle$, so $\langle I\rangle$ is closed under addition. Given $a/q\in\mathbbm{A}_{I}$ and $b/s\in\langle I\rangle$ with $a\in\mathbbm{A}$, $b\in I$, and $q,s\in R\setminus I$, then $(a/q)(b/s)=(ab)/(qs)\in\langle I\rangle$, so $\langle I\rangle$ is a left ideal of $\mathbbm{A}_{I}$.
The inclusion $\subseteq$” in (4) holds by definition of $\langle I\rangle$. Conversely, suppose $a\in\mathbbm{A}$ fulfils $a/1\in\langle I\rangle$. This means there are $b\in I$ and $s\in R\setminus I$ such that $a/1=b/s$; i.e. $sa-b\in[I:\mathbbm{A}]$. Since $[I:\mathbbm{A}]\subseteq I$ and $b\in I$, it follows that $sa\in I$ and so $s\mathbbm{A}a=\mathbbm{A}sa\subseteq I$. Therefore $a\in I$ because $I$ is prime and $s\notin I$. ∎
Recall that any nondegenerate bilinear form $\beta\colon V\times V\to\mathbbm{F}$ on a finite-dimensional $\mathbbm{F}$-vector space $V$ induces an $\mathbbm{F}$-linear isomorphism between $V$ and its dual vector space $V^{*}$, namely $\ignorespaces{\,\cdot\,}\ignorespaces^{\flat}\colon V\to V^{*}$, $w\mapsto w^{\flat}$ with $w^{\flat}(v)\coloneqq\beta(v,w)$ for all $v,w\in V$. In this case it follows that for every linear subspace $U$ of $V$ and any $v\in V\setminus U$ there exists $w\in V$ such that $\beta(u,w)=0$ for all $u\in U$ and $\beta(v,w)\neq 0$.
Theorem 14.
Let $\mathbbm{A}$ be a ring that is finitely generated as a module over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$ and $I$ a prime left ideal of $\mathbbm{A}$. Write
(5) $I^{\perp}\coloneqq{}$$\bigl{\{}b\in\mathbbm{A}\mid Ib\subseteq[I:\mathbbm{A}]\bigr{\}}.$
Then $I^{\perp}$ is a right ideal of $\mathbbm{A}$ and
(6) $I={}$$\bigl{\{}a\in\mathbbm{A}\mid aI^{\perp}\subseteq[I:\mathbbm{A}]\bigr{\}}.$
Proof.
It is easy to check that $I^{\perp}$ is a right ideal of $\mathbbm{A}$ and that the inclusion $\subseteq$” holds in (6), so it only remains to prove the converse inclusion $\supseteq$”. Assume $c\in\mathbbm{A}\setminus I$, then in order to complete the proof we have to find $b\in I^{\perp}$ such that $cb\notin[I:\mathbbm{A}]$. Note that $I\neq\mathbbm{A}$.
By Definition 11 and Remark 12, $\mathbbm{A}_{I}$ is a prime $R_{I}$-algebra. As $\mathbbm{A}$ is finitely generated as an $R$-module by assumption, $\mathbbm{A}_{I}$ has finite dimension over $R_{I}$, i.e. $\mathbbm{A}_{I}$ is a finite-dimensional prime algebra over the field $R_{I}$. By [15, Prop. 11.7] this implies that $\mathbbm{A}_{I}$ is simple, so $\mathbbm{A}_{I}$ is central simple over its center $\mathcal{Z}(\mathbbm{A}_{I})\supseteq R_{I}$. By [16, Lemma 1.1] and [8, Lemma 4 and formulas (9) and (11) in §$\S$22], the reduced trace $\operatorname{Tr}\colon\mathbbm{A}_{I}\to\mathcal{Z}(\mathbbm{A}_{I})$ is a nondegenerate trace functional; i.e. $\operatorname{Tr}$ is a $\mathcal{Z}(\mathbbm{A}_{I})$-linear functional that fulfils $\operatorname{Tr}(xy)=\operatorname{Tr}(yx)$ for all $x,y\in\mathbbm{A}_{I}$, and for every $x\in\mathbbm{A}_{I}\setminus\{0/1\}$ there exists $y\in\mathbbm{A}_{I}$ such that $\operatorname{Tr}(xy)\neq 0/1$.
As $c\in\mathbbm{A}\setminus I$ it follows from (4) of the previous Proposition 13 that $c/1\in\mathbbm{A}_{I}\setminus\langle I\rangle$. As $\langle I\rangle$ is a left ideal of $\mathbbm{A}_{I}$ it is in particular a $\mathcal{Z}(\mathbbm{A}_{I})$-linear subspace. As the bilinear functional $\mathbbm{A}_{I}\times\mathbbm{A}_{I}\ni(x,y)\mapsto\operatorname{Tr}(xy)\in\mathcal{Z}(\mathbbm{A}_{I})$ is nondegenerate and as $\mathbbm{A}_{I}$ is finite-dimensional over $R_{I}$, hence also over $\mathcal{Z}(\mathbbm{A}_{I})$, there exists $y\in\mathbbm{A}_{I}$ such that $\operatorname{Tr}(xy)=0/1$ for all $x\in\langle I\rangle$ and $\operatorname{Tr}\bigl{(}(c/1)y\bigr{)}\neq 0/1$. The former means that $xy=0/1$ for all $x\in\langle I\rangle$, because $\operatorname{Tr}$ is nondegenerate and $\operatorname{Tr}(xyz)=\operatorname{Tr}(zxy)=\operatorname{Tr}(x^{\prime}y)=0/1$ for all $x\in\langle I\rangle$ and all $z\in\mathbbm{A}_{I}$ with $x^{\prime}\coloneqq zx\in\langle I\rangle$. The latter implies that $(c/1)y\neq 0/1$. By definition of $\mathbbm{A}_{I}$ there are $b\in\mathbbm{A}$ and $s\in R\setminus I$ such that $y=b/s$. We have shown that $(ab)/(qs)=(a/q)(b/s)=0/1$ for all $a\in I$, $q\in R\setminus I$, and that $(cb)/s=(c/1)(b/s)\neq 0/1$. The former means that $ab\in[I:\mathbbm{A}]$ for all $a\in I$, i.e. $b\in I^{\perp}$, and the latter that $cb\notin[I:\mathbbm{A}]$. ∎
Corollary 15.
Let $\mathbbm{A}$ be a ring that is finitely generated as a module over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$. Then $\mathbbm{A}$ is weakly left Jacobson if and only if $\mathbbm{A}$ is Jacobson.
Proof.
If $\mathbbm{A}$ is weakly left Jacobson then $\mathbbm{A}$ clearly is Jacobson. Conversely, assume that $\mathbbm{A}$ is Jacobson and let $I$ be a prime left ideal of $\mathbbm{A}$. We show that $I$ is the intersection of all maximal left ideals of $\mathbbm{A}$ that contain $I$. In order to do so consider any $c\in\mathbbm{A}\setminus I$. The proof is complete once we have found a maximal left ideal $\mathfrak{m}$ of $\mathbbm{A}$ that fulfils $I\subseteq\mathfrak{m}$ and $c\notin\mathfrak{m}$.
By the previous Theorem 14 there exists $b\in I^{\perp}$ such that $cb\notin[I:\mathbbm{A}]$. As $[I:\mathbbm{A}]$ is a prime ideal of $\mathbbm{A}$ by Proposition 10, and as $\mathbbm{A}$ is Jacobson by assumption, there exists a maximal left ideal $\tilde{\mathfrak{m}}$ of $\mathbbm{A}$ that fulfils $[I:\mathbbm{A}]\subseteq\tilde{\mathfrak{m}}$ and $cb\notin\tilde{\mathfrak{m}}$. We construct the simple $\mathbbm{A}$-left module $M\coloneqq\mathbbm{A}/\tilde{\mathfrak{m}}$ and write $[a]\in M$ for the equivalence class of an element $a\in\mathbbm{A}$. Then $a[b]=[ab]=[0]$ for all $a\in I$ because $b\in I^{\bot}$ so that $ab\in[I:\mathbbm{A}]\subseteq\tilde{\mathfrak{m}}$, and $c[b]=[cb]\neq[0]$. Note that this also implies that $[b]\neq[0]$. So the annihilator $\mathfrak{m}\coloneqq\operatorname{ann}_{M}([b])=\{a\in\mathbbm{A}\mid a[b]=[0]\}$ is a maximal left ideal of $\mathbbm{A}$ because $M$ is simple (see Lemma 3), and $I\subseteq\mathfrak{m}$ and $c\notin\mathfrak{m}$ hold as required. ∎
Corollary 16.
Let $\mathbbm{A}$ be a ring that is finitely generated as a module over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$. If $R$ is Noetherian, then $\mathbbm{A}$ is weakly left Jacobson.
Proof.
If $R$ is Noetherian, then $\mathbbm{A}$ is Jacobson by [17, Sec. 9.1.3], so $\mathbbm{A}$ is weakly left Jacobson by the previous Corollary 15. ∎
5. Some strongly Jacobson rings
In this section we show that an Azumaya $R$-algebra is strongly left Jacobson if and only if $R$ is Jacobson. This follows from Corollary 15 and Proposition 19; see Theorem 20.
Let the ring $\mathbbm{A}$ be a separable algebra over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$; i.e. $\mathbbm{A}$ is an Azumaya $R$-algebra. By [7, Thm. 3.4 of Chap. II] or [10, Thm. 7.1.4], $\mathbbm{A}$ is a finitely generated projective $R$-module and the morphism of $R$-algebras
(7) $\Xi_{\mathbbm{A}}\colon\mathbbm{A}\otimes_{R}\mathbbm{A}^{\mathrm{op}}\to\mathrm{End}_{R}(\mathbbm{A}),\quad a\otimes b\mapsto\Xi_{\mathbbm{A}}(a\otimes b)$
defined by
(8) $\Xi_{\mathbbm{A}}(a\otimes b)(c)\coloneqq acb\quad\quad\text{for $a,b,c\in\mathbbm{A}$}$
is an isomorphism. Here $\mathbbm{A}^{\mathrm{op}}$ denoted the opposite ring of $\mathbbm{A}$ and $\mathrm{End}_{R}(\mathbbm{A})$ is the $R$-algebra of $R$-module endomorphisms on $\mathbbm{A}$.
Lemma 17.
Let $\mathbbm{A}$ be an Azumaya algebra with center $R\coloneqq\mathcal{Z}(\mathbbm{A})$ and let $(v_{i},\omega_{i})_{i=1,\ldots,d}$ be a dual basis of $M\coloneqq\mathbbm{A}$ considered as a (finitely generated projective) $R$-module; i.e. $v_{1},\dots,v_{d}\in M$ and $\omega_{1},\dots,\omega_{d}\in\operatorname{Hom}_{R}(M,R)$ are such that $\sum_{i=1}^{d}\omega_{i}(x)v_{i}=x$ for every $x\in M$. Then there exist $\rho\in\mathbbm{N}$ and $e_{i}^{\alpha},f_{i}^{\alpha}\in\mathbbm{A}$ for $i\in\{1,\dots,d\}$, $\alpha\in\{1,\dots,\rho\}$ such that for all $i\in\{1,\dots,d\}$ and $x\in\mathbbm{A}$ we have
(9) $\sum_{\alpha=1}^{\rho}e_{i}^{\alpha}xf_{i}^{\alpha}=\omega_{i}(x)$
Proof.
Every functional $\omega_{i}\colon\mathbbm{A}\to R\subseteq\mathbbm{A}$ with $i\in\{1,\dots,d\}$ describes an element of $\mathrm{End}_{R}(\mathbbm{A})$. Since $\Xi_{\mathbbm{A}}\colon\mathbbm{A}\otimes_{R}\mathbbm{A}^{\mathrm{op}}\to\mathrm{End}_{R}(\mathbbm{A})$ is surjective, the preimages of $\omega_{i}$ under $\Xi_{\mathbbm{A}}$ provide elements $e_{i}^{\alpha},f_{i}^{\alpha}\in\mathbbm{A}$ for $i\in\{1,\dots,d\}$ and $\alpha\in\{1,\dots,\rho\}$, with $\rho\in\mathbbm{N}$ sufficiently large, such that (9) holds. ∎
A submodule $N$ of a module $M$ over a commutative ring $R$ is called prime if the following holds: Whenever elements $r\in R$ and $m\in M$ fulfil $rm\in N$, then $rM\subseteq N$ or $m\in N$. In [3, Def. 2.1], a submodule $N$ of a module $M$ is called semiprime if the following holds: Whenever an element $m\in M$ fulfils $m\in\sum(N:m)M$ we have $m\in N$; here $(N:m)$ denotes the ideal $(N:m)\coloneqq\{r\in R\mid rm\in N\}$ of $R$. By [3, Thm. 2.6], every semiprime submodule of a finitely generated $R$-module is an intersection of prime submodules.
Lemma 18.
Suppose $M$ is a finitely generated projective $R$-module. Pick a dual basis $(v_{i},\omega_{i})_{i=1,\ldots,d}$ of $M$. A submodule $N$ of $M$ is semiprime if and only if for every $x\in M$ that fulfils $\omega_{i}(x)\in(N:x)$ for all $i\in\{1,\dots,d\}$ we have $x\in N$.
Proof.
Suppose that $N$ is a semiprime submodule of $M$. Pick $x\in M$ such that $\omega_{i}(x)\in(N:x)$ for all $i\in\{1,\dots,d\}$. Then it follows that $x=\sum_{i=1}^{d}\omega_{i}(x)v_{i}\in\sum(N:x)M$. Since $N$ is semiprime, this implies $x\in N$.
Conversely, consider a submodule $N$ of $M$ such that for any $x\in M$ it follows from $\omega_{i}(x)\in(N:x)$ for all $i\in\{1,\dots,d\}$ that $x\in N$. To prove that $N$ is semiprime, take any $x=\sum_{j=1}^{k}r_{j}m_{j}\in M$ with $k\in\mathbbm{N}_{0}$ and $r_{j}\in(N:x)$, $m_{j}\in M$ for $j\in\{1,\ldots,k\}$. Since $(N:x)$ is an ideal in $R$ we have $\omega_{i}(x)=\sum_{j=1}^{k}r_{j}\omega_{i}(m_{j})\in(N:x)$ for all $i\in\{1,\dots,d\}$, so $x\in N$ by assumption. ∎
Proposition 19.
Let the ring $\mathbbm{A}$ be an Azumaya algebra over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$. Then $\mathbbm{A}$ is a left intersection ring, i.e. every semiprime left ideal of $\mathbbm{A}$ is an intersection of prime left ideals of $\mathbbm{A}$.
Proof.
Let $I$ be a semiprime left ideal of $\mathbbm{A}$ and consider an arbitrary element $c\in\mathbbm{A}\setminus I$. In order to complete the proof we construct a prime left ideal $J$ of $\mathbbm{A}$ that fulfils $I\subseteq J$ and $c\notin J$.
Let $(v_{i},\omega_{i})_{i=1,\ldots,d}$ be a dual basis of $\mathbbm{A}$ considered as an $R$-module. The first step is to show that $I$ is also semiprime as a submodule of $\mathbbm{A}$. So assume that $x\in\mathbbm{A}$ satisfies $\omega_{i}(x)\in(I:x)$ for all $i\in\{1,\ldots,d\}$. Note that $\omega_{i}(x)\in(I:x)$ means that $\omega_{i}(x)x\in I$. Then $xax=\sum_{i=1}^{d}\omega_{i}(x)v_{i}ax=\sum_{i=1}^{d}v_{i}a\,\omega_{i}(x)x\in I$ for all $a\in\mathbbm{A}$, which implies $x\in I$ because $I$ is a semiprime left ideal of $\mathbbm{A}$. By Lemma 18 this means that $I$ is indeed a semiprime submodule of $\mathbbm{A}$.
By [3, Thm. 2.6] there exists a prime $R$-submodule $N$ of $\mathbbm{A}$ such that $I\subseteq N$ and $c\notin N$. Set $J\coloneqq\{x\in\mathbbm{A}\mid\mathbbm{A}x\subseteq N\}$. It is easy to check that $J$ is a left ideal of $N$ and $J\subseteq N$ because $\mathbbm{A}$ has a unit. In particular $c\notin J$. Moreover, $I\subseteq J$ because $\mathbbm{A}x\subseteq I\subseteq N$ for all $x\in I$. It only remains to show that $J$ is prime as a left ideal of $\mathbbm{A}$:
Let $\rho\in\mathbbm{N}$ and $e_{i}^{\alpha},f_{i}^{\alpha}\in\mathbbm{A}$ for $i\in\{1,\dots,d\}$, $\alpha\in\{1,\dots,\rho\}$ fulfil formula (9) of Lemma 17. Pick any $x\in\mathbbm{A}$ and $x^{\prime}\in\mathbbm{A}\setminus J$ such that $xax^{\prime}\in J$ for all $a\in\mathbbm{A}$. As $x^{\prime}\notin J$ there is $a^{\prime}\in\mathbbm{A}$ such that $a^{\prime}x^{\prime}\notin N$, On the other hand, for each $i\in\{1,\ldots,d\}$ we have $\omega_{i}(x)a^{\prime}x^{\prime}=\sum_{\alpha=1}^{\rho}e_{i}^{\alpha}xf_{i}^{\alpha}a^{\prime}x^{\prime}\in J\subseteq N$. Since $N$ is a prime $R$-submodule of $\mathbbm{A}$ and $a^{\prime}x^{\prime}\notin N$ we have $\omega_{i}(x)\mathbbm{A}\subseteq N$ for all $i\in\{1,\ldots,d\}$. It follows that $ax=a\sum_{i=1}^{d}\omega_{i}(x)v_{i}=\sum_{i=1}^{d}\omega_{i}(x)av_{i}\in N$ for all $a\in\mathbbm{A}$, so $x\in J$. ∎
By [7, Cor. 3.7 of Chap. II] or [10, Cor. 7.1.7] there is a bijective correspondence between the (two-sided) ideals of an Azumaya algebra $\mathbbm{A}$ and the ideals of its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$. It is given by
(10) $I\mapsto I\cap R\quad\quad\text{and}\quad\quad J\mapsto{\textstyle\sum}J\mathbbm{A}$
for (two-sided) ideals $I$ of $\mathbbm{A}$ and ideals $J$ or $R$, respectively. It is clear that the relation $\subseteq$ is preserved in both directions, and consequently also maximal ideals are preserved. As observed in [11, Thm. 8.17(b)] the correspondence also preserves prime ideals.
Theorem 20.
Let $\mathbbm{A}$ be an Azumaya algebra over its center $R\coloneqq\mathcal{Z}(\mathbbm{A})$, then the following holds:
  1. i)
    $\mathbbm{A}$ is strongly left Jacobson if and only if $R$ is Jacobson.
  2. ii)
    If $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}$, then the (two-sided) ideal $[\mathfrak{m}:\mathbbm{A}]$ of $\mathbbm{A}$ from (2) is maximal, $[\mathfrak{m}:\mathbbm{A}]\cap R=\mathfrak{m}\cap R$, and $\mathfrak{m}\cap R$ is a maximal ideal of $R$.
Proof.
We begin with part ii): Let $\mathfrak{m}$ be a maximal left ideal of $\mathbbm{A}$. Note that $[\mathfrak{m}:\mathbbm{A}]$ is the annihilator of the simple left $\mathbbm{A}$-module $\mathbbm{A}/\mathfrak{m}$, so $[\mathfrak{m}:\mathbbm{A}]$ is left-primitive. Since $\mathbbm{A}$ is a PI ring by [11, Lemma 1.19], every left-primitive ideal of $\mathbbm{A}$ is maximal by Kaplansky’s Theorem [11, Thm. 4.13]; in particular $[\mathfrak{m}:\mathbbm{A}]$ is a maximal ideal of $\mathbbm{A}$. Since the correspondence (10) preserves maximal ideals, $[\mathfrak{m}:\mathbbm{A}]\cap R$ is a maximal ideal of $R$. By Proposition 10, $[\mathfrak{m}:\mathbbm{A}]\cap R=\mathfrak{m}\cap R$.
We proceed with part i). By the correspondence (10), $R$ is Jacobson (i.e. every prime ideal of $R$ is an intersection of maximal ideals of $R$) if and only if every prime ideal of $\mathbbm{A}$ is an intersection of maximal ideals of $\mathbbm{A}$. The latter in turn is true if and only if every prime ideal of $\mathbbm{A}$ is an intersection of left-primitive ideals (because the left-primitive ideals in $\mathbbm{A}$ coincide with the maximal ideals by the previous paragraph and [15, Prop. 11.6]). Recall that, by [15, Cor. 11.5], $\mathbbm{A}$ is Jacobson if and only if every prime ideal of $\mathbbm{A}$ is an intersection of left-primitive ideals, so we have shown that $R$ is Jacobson if and only if $\mathbbm{A}$ is Jacobson. By Corollary 15 and Proposition 19, $\mathbbm{A}$ is Jacobson if and only if $\mathbbm{A}$ is strongly left Jacobson. ∎
Remark 21.
Part ii) of the above Theorem 20 implies that for every maximal left ideal $\mathfrak{m}$ of $\mathbbm{A}$ the simple left module $\mathbbm{A}/\mathfrak{m}$ is an $R/(\mathfrak{m}\cap R)$-vector space. As $\mathbbm{A}$ is finitely generated over $R$, the dimension of $\mathbbm{A}/\mathfrak{m}$ over $R/(\mathfrak{m}\cap R)$ is finite.
6. Polynomials with coefficients in algebras
The aim of this section is to show that every ring of the form $\mathbbm{A}[x_{1},\ldots,x_{n}]\coloneqq\mathbbm{A}\otimes_{\mathbbm{F}}\mathbbm{F}[x_{1},\dots,x_{n}]$, where $\mathbbm{A}$ is a finite-dimensional algebra over a field $\mathbbm{F}$ and $x_{1},\ldots,x_{n}$ are central variables, is strongly left Jacobson; see Theorem 27. The cases $\mathbbm{A}=\mathbb{H}$ and $\mathbbm{A}=M_{k}(\mathbb{C})$ were already proved in [6, Theorems B and D]. The proof of Theorem 27 consists of three steps. For simple $\mathbbm{A}$ we use Theorem 20, then we prove it for semisimple $\mathbbm{A}$ and finally for general $\mathbbm{A}$. In parallel we also prove that every maximal left ideal of $\mathbbm{A}[x_{1},\ldots,x_{n}]$ has a special form, see Proposition 28, which implies that every semiprime left ideal is equal to the intersection of all ideals of the special form that contain it; see Corollary 29. This result can be interpreted as a one-sided Nullstellensatz for $\mathbbm{A}[x_{1},\ldots,x_{n}]$ in the spirit of [6, Theorems B and D].
If $\mathbbm{A}$ is a finite-dimensional and simple $\mathbbm{F}$-algebra, then its center $\mathbbm{E}\coloneqq\mathcal{Z}(\mathbbm{A})$ is a finite (hence algebraic) field extension of $\mathbbm{F}$. Write $\bar{\mathbbm{F}}$ for the algebraic closure of $\mathbbm{F}$, then $\bar{\mathbbm{F}}$ is also the algebraic closure of $\mathbbm{E}$, so by [10, Prop. 4.5.1.], there exists $k\in\mathbbm{N}$ such that $\mathbbm{A}\otimes_{\mathbbm{E}}\bar{\mathbbm{F}}\cong\mathrm{M}_{k}(\bar{\mathbbm{F}})$ where $\mathrm{M}_{k}(\bar{\mathbbm{F}})$ denotes the $\bar{\mathbbm{F}}$-algebra of square matrices of size $k$.
For $\xi\in\bar{\mathbbm{F}}^{n}$ an evaluation map $\mathbbm{A}[x_{1},\dots,x_{n}]\ni a\mapsto a(\xi)\in\mathrm{M}_{k}(\bar{\mathbbm{F}})$ can be defined by
(11) $a(\xi)\coloneqq\sum_{\nu}a_{\nu}\xi^{\nu}\in\mathbbm{A}\otimes_{\mathbbm{E}}\bar{\mathbbm{F}}\cong\mathrm{M}_{k}(\bar{\mathbbm{F}})$
for all $a=\sum_{\nu}a_{\nu}x^{\nu}\in\mathbbm{A}[x_{1},\dots,x_{n}]$ with coefficients $a_{\nu}\in\mathbbm{A}$ (using standard multiindex notation). This evaluation map clearly is an $\mathbbm{F}$-algebra morphism, and therefore
(12) $J_{\xi,v}\coloneqq\bigl{\{}a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)v=0\bigr{\}}$
is a left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ for every $v\in\bar{\mathbbm{F}}^{k}$. Note that $1\in J_{\xi,v}$ if and only if $v=0$.
Proposition 22.
Let $\mathbbm{F}$ be a field, $\mathbbm{A}$ a finite-dimensional simple $\mathbbm{F}$-algebra, and $n\in\mathbbm{N}$. Then $\mathbbm{A}[x_{1},\dots,x_{n}]$ is strongly left Jacobson. Moreover, if $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, then the (two-sided) ideal $\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ from (2) is maximal, $\mathfrak{m}$ has finite codimension over $\mathbbm{F}$, and there exist a point $\xi\in\bar{\mathbbm{F}}^{n}$ and a vector $v\in\bar{\mathbbm{F}}^{k}\setminus\{0\}$ such that $\mathfrak{m}=J_{\xi,v}$.
Proof.
Let $\mathbbm{E}$ be the center of $\mathbbm{A}$ so that the center of $\mathbbm{A}[x_{1},\dots,x_{n}]$ is $R\coloneqq\mathcal{Z}\bigl{(}\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{)}=\mathbbm{E}[x_{1},\ldots,x_{n}]$. Since $\mathbbm{A}$ is central simple over $\mathbbm{E}$ it is also Azumaya by [10, Cor. 4.5.4.]. By [10, Cor. 4.3.2] it follows that $\mathbbm{A}[x_{1},\dots,x_{n}]\cong\mathbbm{A}\otimes_{\mathbbm{E}}R$ is Azumaya over $R$. We can now apply Theorem 20. By Hilbert’s Nullstellensatz, the polynomial ring $R$ is Jacobson (see e.g. [9, Thm. 4.9] or [12, Cor. on p. 138]) and every maximal ideal of $R$ has finite codimension over $\mathbbm{E}$ (see e.g [4, Cor. 7.10] or [12, Cor. on p. 139]). Part i) of Theorem 20 now shows that $\mathbbm{A}[x_{1},\dots,x_{n}]$ is strongly left Jacobson.
Moreover, if $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, then by part ii) of Theorem 20, $\mathfrak{n}\coloneqq\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}$ is a maximal (two-sided) ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, and $\mathfrak{n}\cap R=\mathfrak{m}\cap R$ is a maximal ideal of $R$ and in particular has finite codimension over $\mathbbm{E}$ in $R$. Write $\mathbbm{K}\coloneqq R/(\mathfrak{m}\cap R)$, then $\mathbbm{K}$ is a finite field extension of $\mathbbm{E}$. By Remark 21, $\mathbbm{A}[x_{1},\dots,x_{n}]/\mathfrak{m}$ is finite dimensional over $\mathbbm{K}$ and therefore
$\dim_{\mathbbm{F}}\mathbbm{A}[x_{1},\dots,x_{n}]/\mathfrak{m}=\dim_{\mathbbm{F}}\mathbbm{E}\cdot\dim_{\mathbbm{E}}\mathbbm{K}\cdot\dim_{\mathbbm{K}}\mathbbm{A}[x_{1},\dots,x_{n}]/\mathfrak{m}<\infty$
holds. Moreover, as $\mathbbm{K}$ is finite dimensional (hence algebraic) also over $\mathbbm{F}$, it embedds into the algebraic closure $\bar{\mathbbm{F}}$. Set $\xi\coloneqq(\xi_{1},\dots,\xi_{n})\coloneqq\bigl{(}[x_{1}],\dots,[x_{n}]\bigr{)}\in\mathbbm{K}^{n}\subseteq\bar{\mathbbm{F}}^{n}$ where $[\ignorespaces{\,\cdot\,}\ignorespaces]\colon R\to\mathbbm{K}$ is the canonical projection onto the quotient, which is an $\mathbbm{E}$-algebra morphism. Then $\mathfrak{m}\cap R=\ker[\ignorespaces{\,\cdot\,}\ignorespaces]=\{p\in R\mid p(\xi)=0\}$ because $p(\xi)=[p(x_{1},\dots,x_{n})]=[p]$ for all $p\in R$.
We claim that $\mathfrak{n}=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)=0\}$: As $\mathbbm{A}[x_{1},\dots,x_{n}]$ is Azumaya over $R$ we have the bijective correspondence (10) between the (two-sided) ideals of $\mathbbm{A}[x_{1},\dots,x_{n}]$ and the ideals of $R$, so $\mathfrak{n}=\sum(\mathfrak{m}\cap R)\mathbbm{A}[x_{1},\dots,x_{n}]$ because we already know that $\mathfrak{n}\cap R=\mathfrak{m}\cap R$. From this it follows that $\mathfrak{n}\subseteq\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)=0\}$. As evaluation at $\xi$ is an $\mathbbm{E}$-algebra morphism $\mathbbm{A}[x_{1},\dots,x_{n}]\to\mathrm{M}_{k}(\bar{\mathbbm{F}})$, its kernel $\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)=0\}$ is a proper (two-sided) ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, so $\mathfrak{n}=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)=0\}$ by maximality of $\mathfrak{n}$.
Write $\mathfrak{m}^{\bot}\coloneqq\{b\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid\mathfrak{m}\,b\subseteq\mathfrak{n}\}$ like in Theorem 14, then $\mathfrak{m}=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a\,\mathfrak{m}^{\bot}\subseteq\mathfrak{n}\}$, or equivalently,
$\mathfrak{m}=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a(\xi)\,b(\xi)=0\leavevmode\nobreak\ \textup{for all $b\in\mathfrak{m}^{\bot}$}\}.$
As $1\notin\mathfrak{m}$ there is $b\in\mathfrak{m}^{\bot}$ such that $b(\xi)\neq 0$. Let $v\in\bar{\mathbbm{F}}^{k}\setminus\{0\}$ be any non-zero column of $b(\xi)\in\mathrm{M}_{k}(\bar{\mathbbm{F}})$, then $a(\xi)v=0$ for all $a\in\mathbbm{A}[x_{1},\dots,x_{n}]$, so $\mathfrak{m}\subseteq J_{\xi,v}$. By maximality of $\mathfrak{m}$ it follows that $\mathfrak{m}=J_{\xi,v}$. ∎
Example 23.
Note, however, that not every left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ of the form $J_{\xi,v}$ is maximal. Namely, pick $\mathbbm{A}=M_{2}(\mathbbm{Q})$, $\xi=(0,\ldots,0)\in\mathbbm{Q}^{n}$ and $v=[1\ \sqrt{2}]^{T}\in\bar{\mathbbm{Q}}^{2}$. Since $\sqrt{2}$ is irrational, $a(\xi)v=0$ is equivalent to $a(\xi)=0$ for every $a\in\mathbbm{A}[x_{1},\ldots,x_{n}]$. To show that $J_{\xi,v}$ is not a maximal left ideal pick any $0\neq u\in\mathbbm{Q}^{2}$ and note that $a(\xi)u=0$ need not imply $a(\xi)=0$. Therefore $J_{\xi,v}\subsetneq J_{\xi,u}\subsetneq\mathbbm{A}[x_{1},\dots,x_{n}]$.
Proposition 24.
Consider rings $\mathbbm{A}_{1},\ldots,\mathbbm{A}_{k}$ with $k\in\mathbbm{N}$ and let $\mathbbm{A}\coloneqq\mathbbm{A}_{1}\oplus\dots\oplus\mathbbm{A}_{k}$. For any $j\in\{1,\dots,k\}$ and any left ideal $I_{j}$ of $\mathbbm{A}_{j}$ we write
(13) $\hat{I}_{j}\coloneqq\mathbbm{A}_{1}\oplus\dots\oplus\mathbbm{A}_{j-1}\oplus I_{j}\oplus\mathbbm{A}_{j+1}\oplus\dots\oplus\mathbbm{A}_{k}\subseteq\mathbbm{A}.$
Then a left ideal $\mathfrak{m}$ of $\mathbbm{A}$ is maximal if and only if there exist $j\in\{1,\dots,k\}$ and a maximal ideal $\mathfrak{m}_{j}$ of $\mathbbm{A}_{j}$ such that $\mathfrak{m}=\hat{\mathfrak{m}}_{j}$. Moreover, if all $\mathbbm{A}_{1},\dots,\mathbbm{A}_{k}$ are strongly left Jacobson, then $\mathbbm{A}$ is strongly left Jacobson.
Proof.
Write $p_{j}\coloneqq 0\oplus\dots\oplus 0\oplus 1\oplus 0\oplus\dots\oplus 0\in\mathbbm{A}_{j}\subseteq\mathbbm{A}$ for $j\in\{1,\dots,k\}$, where $1$ denotes the unit of $\mathbbm{A}_{j}$ and is situated at position $j$. Clearly $\sum_{j=1}^{k}p_{j}$ is the unit of $\mathbbm{A}$. Every left ideal $I$ of $\mathbbm{A}$ decomposes as
$I=I_{1}\oplus\dots\oplus I_{k}=\hat{I}_{1}\cap\dots\cap\hat{I}_{k},$
where, for each $j\in\{1,\ldots,k\}$, $I_{j}\coloneqq p_{j}I=I\cap\mathbbm{A}_{j}$ is a left ideal of $A_{j}$ and $\hat{I}_{j}$ is a left ideal of $\mathbbm{A}$. In particular if $I=\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}$, then by maximality of $\mathfrak{m}$, $\hat{\mathfrak{m}}_{j}=\mathbbm{A}$ for all but one $j\in\{1,\dots,k\}$, so $\mathfrak{m}=\hat{\mathfrak{m}}_{j}$ for this remaining $j\in\{1,\dots,k\}$. Conversely, if $\mathfrak{n}_{j}$ is any maximal left ideal of $\mathbbm{A}_{j}$ for some $j\in\{1,\dots,k\}$, then it is clear that $\hat{\mathfrak{n}}_{j}$ is a maximal left ideal of $\mathbbm{A}$.
Let $I$ be a semiprime left ideal of $\mathbbm{A}$. For any $j\in\{1,\dots,k\}$, the left ideal $I_{j}$ of $\mathbbm{A}_{j}$ is also semiprime: indeed, if $a\in\mathbbm{A}_{j}$ fulfils $a\mathbbm{A}_{j}a\subseteq I_{j}$, then $a\mathbbm{A}a=ap_{j}\mathbbm{A}p_{j}a\subseteq I_{j}\subseteq I$ because $p_{j}\mathbbm{A}p_{j}\subseteq\mathbbm{A}_{j}$, so $a\in I$ because $I$ is semiprime, and therefore $a\in I_{j}$. Now consider any element $\hat{c}=c_{1}\oplus\dots\oplus c_{k}\in\mathbbm{A}\setminus I$ with components $c_{j}\in\mathbbm{A}_{j}$ for $j\in\{1,\dots,k\}$. Then there exists $j\in\{1,\dots,k\}$ such that $c_{j}\in\mathbbm{A}_{j}\setminus I_{j}$, and if $\mathbbm{A}_{j}$ is strongly left Jacobson, then there is a maximal left ideal $\mathfrak{m}_{j}$ of $\mathbbm{A}_{j}$ such that $I_{j}\subseteq\mathfrak{m}_{j}$ and $c_{j}\notin\mathfrak{m}_{j}$. In this case $\hat{\mathfrak{m}}_{j}$ is a maximal left ideal of $\mathbbm{A}$ that fulfils $I\subseteq\hat{\mathfrak{m}}_{j}$ and $\hat{c}\notin\hat{\mathfrak{m}}_{j}$. ∎
Corollary 25.
Let $\mathbbm{F}$ be a field, $\mathbbm{A}$ a finite-dimensional semisimple $\mathbbm{F}$-algebra, and $n\in\mathbbm{N}$. Then $\mathbbm{A}[x_{1},\dots,x_{n}]$ is strongly left Jacobson. Moreover, if $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, then $\mathfrak{m}$ has finite codimension over $\mathbbm{F}$ and the (two-sided) ideal $\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ from (2) is maximal.
Proof.
As $\mathbbm{A}$ is semisimple there are $k\in\mathbbm{N}$ and simple $\mathbbm{F}$-algebras $\mathbbm{A}_{1},\dots,\mathbbm{A}_{k}$ such that $\mathbbm{A}\cong\mathbbm{A}_{1}\oplus\dots\oplus\mathbbm{A}_{k}$. As $\otimes$ is distributive over $\oplus$, $\mathbbm{A}[x_{1},\dots,x_{n}]\cong\mathbbm{A}_{1}[x_{1},\dots,x_{n}]\oplus\dots\oplus\mathbbm{A}_{k}[x_{1},\dots,x_{m}]$. By Proposition 22 each $\mathbbm{F}$-algebra $\mathbbm{A}_{j}[x_{1},\dots,x_{n}]$ with $j\in\{1,\dots,k\}$ is strongly left Jacobson, every maximal left ideal $\mathfrak{m}_{j}$ of $\mathbbm{A}_{j}[x_{1},\dots,x_{n}]$ has finite codimension, and $\mathfrak{n}_{j}\coloneqq\bigl{[}\mathfrak{m}_{j}:\mathbbm{A}_{j}[x_{1},\dots,x_{n}]\bigr{]}$ is a maximal (two-sided) ideal of $\mathbbm{A}_{j}[x_{1},\dots,x_{n}]$. The previous Proposition 24 now shows that $\mathbbm{A}[x_{1},\dots,x_{n}]$ itself is strongly left Jacobson and that every maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ is of the form $\hat{\mathfrak{m}}_{j}$ as in (13) for some $j\in\{1,\dots,m\}$ and some maximal left ideal $\mathfrak{m}_{j}$ of $\mathbbm{A}_{j}[x_{1},\dots,x_{n}]$. Clearly the codimension of $\hat{\mathfrak{m}}_{j}$ in $\mathbbm{A}[x_{1},\dots,x_{n}]$ equals the codimension of $\mathfrak{m}_{j}$ in $\mathbbm{A}_{j}[x_{1},\dots,x_{n}]$, which is finite, and $\bigl{[}\hat{\mathfrak{m}}_{j}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}=\hat{\mathfrak{n}}_{j}$ is a maximal (two-sided) ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$. ∎
We need only one more technical Lemma before we can proof our main theorem:
Lemma 26.
Let $\mathbbm{A}$ be a finite-dimensional algebra over some field $\mathbbm{F}$ and $\mathbbm{B}\coloneqq\mathbbm{A}/\operatorname{rad}\mathbbm{A}$ with $\operatorname{rad}\mathbbm{A}:=\operatorname{rad}(\{0\})$ the Jacobson radical of $\mathbbm{A}$. Write $\tilde{\pi}\colon\mathbbm{A}\to\mathbbm{B}$ for the canonical projection and define
(14) $\pi\coloneqq\tilde{\pi}\otimes\mathrm{id}_{\mathbbm{F}[x_{1},\dots,x_{n}]}\colon\mathbbm{A}[x_{1},\dots,x_{n}]\to\mathbbm{B}[x_{1},\dots,x_{n}].$
Then every semiprime left ideal of $\mathbbm{A}[x_{1},\ldots,x_{n}]$ contains $\ker\pi$.
Proof.
Consider a semiprime left ideal $I$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ and let $J\coloneqq\bigl{[}I\colon\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid a\mathbbm{A}[x_{1},\dots,x_{n}]\subseteq I\}$ be the (two-sided) ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ like in (2). Then $J$ is again semiprime: Indeed, if $a\in\mathbbm{A}[x_{1},\dots,x_{n}]$ fulfils $aca\in J$ for all $c\in\mathbbm{A}[x_{1},\dots,x_{n}]$, then this means that $acad\in I$ for all $c,d\in\mathbbm{A}[x_{1},\dots,x_{n}]$, and therefore also $adcad\in I$ for all $c,d\in\mathbbm{A}[x_{1},\dots,x_{n}]$. But this implies $ad\in I$ for all $d\in\mathbbm{A}[x_{1},\dots,x_{n}]$ because $I$ is semiprime, so $a\in J$.
Next we show that $J\cap\mathbbm{A}$ is a semiprime (two-sided) ideal of $\mathbbm{A}$: It is clear that $J\cap\mathbbm{A}$ is a two-sided ideal of $\mathbbm{A}$. Let $d\in\mathbbm{N}$ be the dimension of $\mathbbm{A}$ over $\mathbbm{F}$ and let $b_{1},\dots,b_{d}$ be an $\mathbbm{F}$-basis of $\mathbbm{A}$. If $a\in\mathbbm{A}$ fulfils $aca\in J\cap\mathbbm{A}$ for all $c\in\mathbbm{A}$, then also $a(\sum_{i=1}^{d}p_{i}b_{i})a=\sum_{i=1}^{d}p_{i}ab_{i}a\in J$ for all elements $\sum_{i=1}^{d}p_{i}b_{i}\in\mathbbm{A}[x_{1},\dots,x_{n}]$ with coefficients $p_{1},\dots,p_{d}\in\mathbbm{F}[x_{1},\dots,x_{n}]$, so $a\in J$ because $J$ is a semiprime ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$.
By [15, Thm. 4.12], $\operatorname{rad}\mathbbm{A}$ is a nilpotent ideal; i.e. $(\operatorname{rad}\mathbbm{A})^{n}=\{0\}\subseteq J\cap\mathbbm{A}$ for some $n$. Since $J\cap\mathbbm{A}$ is semiprime, it follows that $\operatorname{rad}\mathbbm{A}\subseteq J\cap\mathbbm{A}$. Using standard multiindex notation, any element $a\in\mathbbm{A}[x_{1},\ldots,x_{n}]$ can be expanded as a finite sum $a=\sum_{\nu}a_{\nu}x^{\nu}$ with suitable coefficients $a_{\nu}\in\mathbbm{A}$. If $a\in\ker\pi$, then $0=\pi(a)=\sum_{\nu}\tilde{\pi}(a_{\nu})x^{\nu}$, so $\tilde{\pi}(a_{\nu})=0$ for all multiindices $\nu$. Therefore $a_{\nu}\in\operatorname{rad}\mathbbm{A}\subseteq J\cap\mathbbm{A}$ for all multiindices $\nu$, hence $a\in J$. This shows that $\ker\pi\subseteq J\subseteq I$. ∎
Theorem 27.
Let $\mathbbm{A}$ be a finite-dimensional algebra over some field $\mathbbm{F}$ and $n\in\mathbbm{N}$. Then $\mathbbm{A}[x_{1},\dots,x_{n}]$ is strongly left Jacobson. Moreover, if $\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, then $\mathfrak{m}$ has finite codimension over $\mathbbm{F}$ and the (two-sided) ideal $\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ from (2) is maximal.
Proof.
Recall that if we factor a finite-dimensional algebra with its Jacobson radical then we get a semisimple algebra; see [15, Thm. 4.14]. Corollary 25 therefore applies to the algebra $\mathbbm{B}[x_{1},\dots,x_{n}]$ from the previous Lemma 26 and shows that $\mathbbm{B}[x_{1},\dots,x_{n}]$ is strongly left Jacobson, that every maximal left ideal $\tilde{\mathfrak{m}}$ of $\mathbbm{B}[x_{1},\dots,x_{n}]$ has finite codimension over $\mathbbm{F}$ and that $\bigl{[}\tilde{\mathfrak{m}}:\mathbbm{B}[x_{1},\dots,x_{n}]\bigr{]}$ is a maximal (two-sided) ideal of $\mathbbm{B}[x_{1},\dots,x_{n}]$.
As the kernel of the surjective algebra morphism $\pi\colon\mathbbm{A}[x_{1},\dots,x_{n}]\to\mathbbm{B}[x_{1},\dots,x_{n}]$ from Lemma 26 is contained in every semiprime left ideal of $\mathbbm{A}[x_{1},\ldots,x_{n}]$, Lemma 6 shows that the image and preimage under $\pi$ give a bijective correspondence between the semiprime (resp. prime, maximal) left ideals of $\mathbbm{A}[x_{1},\dots,x_{n}]$ and semiprime (resp. prime, maximal) left ideals of $\mathbbm{B}[x_{1},\dots,x_{n}]$. It follows that also $\mathbbm{A}[x_{1},\dots,x_{n}]$ is strongly left Jacobson and that every maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ has finite codimension over $\mathbbm{F}$.
We check that $\bigl{[}I:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}=\pi^{-1}\bigl{(}\bigl{[}\pi(I):\mathbbm{B}[x_{1},\dots,x_{n}]\bigr{]}\bigr{)}$ holds for every semiprime left ideal $I$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$: If an element $a\in\mathbbm{A}[x_{1},\dots,x_{n}]$ fulfils $a\mathbbm{A}[x_{1},\dots,x_{n}]\subseteq I$, then applying $\pi$ shows that $\pi(a)\mathbbm{B}[x_{1},\dots,x_{n}]\subseteq\pi(I)$, which proves the inclusion $\subseteq$”. Conversely, if $a\in\mathbbm{A}[x_{1},\dots,x_{n}]$ fulfils $\pi(a)\in\bigl{[}\pi(I):\mathbbm{B}[x_{1},\dots,x_{n}]\bigr{]}$, i.e. $\pi(a)\mathbbm{B}[x_{1},\dots,x_{n}]\subseteq\pi(I)$, then $\pi\bigl{(}a\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{)}\subseteq\pi(I)$, from which it follows that $a\mathbbm{A}[x_{1},\dots,x_{n}]\subseteq I$ because $\ker\pi\subseteq I$ by Lemma 26.
In particular if $I=\mathfrak{m}$ is a maximal left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$, then $\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}=\pi^{-1}\bigl{(}\bigl{[}\pi(\mathfrak{m}):\mathbbm{B}[x_{1},\dots,x_{n}]\bigr{]}\bigr{)}$ is a maximal (two-sided) ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ because $\bigl{[}\pi(\mathfrak{m}):\mathbbm{B}[x_{1},\dots,x_{n}]\bigr{]}$ is a maximal (two-sided) ideal of $\mathbbm{B}[x_{1},\dots,x_{n}]$. ∎
Finally we want to reformulate the above noncommutative Nullstellensatz in a more explicit way. So let $\mathbbm{A}$ be a finite-dimensional algebra over a field $\mathbbm{F}$ and $n\in\mathbbm{N}$. Let $\theta\colon\mathbbm{A}\to\mathbbm{B}$ be a surjective $\mathbbm{F}$-algebra morphism onto a simple $\mathbbm{F}$-algebra $\mathbbm{B}$, then define
(15) $\Theta\coloneqq\theta\otimes_{\mathbbm{F}}\mathrm{id}_{\mathbbm{F}[x_{1},\dots,x_{n}]}\colon\mathbbm{A}[x_{1},\dots,x_{n}]\to\mathbbm{B}[x_{1},\dots,x_{n}].$
Write $\mathbbm{E}\coloneqq\mathcal{Z}(\mathbbm{B})$ for the center of $\mathbbm{B}$, then like in the discussion at the beginning of this section, $\mathbbm{B}\otimes_{\mathbbm{E}}\bar{\mathbbm{F}}\cong\mathrm{M}_{k}(\bar{\mathbbm{F}})$ for some $k\in\mathbbm{N}$. Let $\xi\in\bar{\mathbbm{F}}^{n}$ and $v\in\bar{\mathbbm{F}}^{k}$, then
(16) $J_{\theta,\xi,v}\coloneqq\bigl{\{}a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid\Theta(a)(\xi)v=0\bigr{\}}$
is a left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$ and $1\in J_{\theta,\xi,v}$ if and only if $v=0$.
Proposition 28.
Let $\mathbbm{A}$ be a finite-dimensional algebra over some field $\mathbbm{F}$ and $n\in\mathbbm{N}$. Then for every maximal left ideal $\mathfrak{m}$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ there exist a surjective $\mathbbm{F}$-algebra morphism $\theta\colon\mathbbm{A}\to\mathbbm{B}$ onto a simple $\mathbbm{F}$-algebra $\mathbbm{B}$, a point $\xi\in\bar{\mathbbm{F}}^{n}$, and a vector $v\in\bar{\mathbbm{F}}^{k}$ such that $\mathfrak{m}=J_{\theta,\xi,v}$.
Proof.
The (two-sided) ideal $\mathfrak{n}\coloneqq\bigl{[}\mathfrak{m}:\mathbbm{A}[x_{1},\dots,x_{n}]\bigr{]}$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ is maximal by the previous Theorem 27. We show that the ideal $\mathfrak{n}\cap\mathbbm{A}$ of $\mathbbm{A}$ is also maximal: Let $a\in\mathbbm{A}\setminus(\mathfrak{n}\cap\mathbbm{A})$ be given, then by maximality of $\mathfrak{n}$ there are $\ell\in\mathbbm{N}$ and $b_{1},\dots,b_{\ell},c_{1},\dots,c_{\ell}\in\mathbbm{A}[x_{1},\dots,x_{n}]$ such that $1-\sum_{j=1}^{\ell}b_{j}ac_{j}\in\mathfrak{n}$. Expanding $b_{j}=\sum_{\nu}b_{j,\nu}x^{\nu}$ and $c_{j}=\sum_{\nu}c_{j,\nu}x^{\nu}$ for all $j\in\{1,\dots,\ell\}$ with coefficients $b_{j,\nu},c_{j,\nu}\in\mathbbm{A}$, we find for the 00-th order in the variables $x_{1},\dots,x_{n}$ that $1-\sum_{j=1}^{\ell}b_{j,0}\,a\,c_{j,0}\in\mathfrak{n}\cap\mathbbm{A}$.
Define $\mathbbm{B}\coloneqq\mathbbm{A}/(\mathfrak{n}\cap\mathbbm{A})$ and let $\theta\colon\mathbbm{A}\to\mathbbm{B}$ be the canonical projection onto the quotient. Then $\mathbbm{B}$ is a simple $\mathbbm{F}$-algebra and $\theta$ a surjective $\mathbbm{F}$-algebra morphism. Define $\Theta\colon\mathbbm{A}[x_{1},\dots,x_{n}]\to\mathbbm{B}[x_{1},\dots,x_{n}]$ as in (15), then an element $a=\sum_{\nu}a_{\nu}x^{\nu}\in\mathbbm{A}[x_{1},\dots,x_{n}]$ with coefficients $a_{\nu}\in\mathbbm{A}$ fulfils $\Theta(a)=0$ if and only of $a_{\nu}\in\mathfrak{n}\cap\mathbbm{A}$ for all multiindices $\nu$, which implies $a\in\mathfrak{n}\subseteq\mathfrak{m}$. So $\ker\Theta\subseteq\mathfrak{m}$. It follows that $\Theta(\mathfrak{m})$ is a maximal left ideal of $\mathbbm{B}[x_{1},\dots,x_{n}]$ and $\mathfrak{m}=\Theta^{-1}(\Theta(\mathfrak{m}))$. By Proposition 22 there are $\xi\in\bar{\mathbbm{F}}^{n}$ and $v\in\bar{\mathbbm{F}}^{k}$ such that $\Theta(\mathfrak{m})=\{b\in\mathbbm{B}[x_{1},\dots,x_{n}]\mid b(\xi)v=0\}$, so $\mathfrak{m}=\Theta^{-1}(\Theta(\mathfrak{m}))=\{a\in\mathbbm{A}[x_{1},\dots,x_{n}]\mid\Theta(a)(\xi)v=0\}$. ∎
Corollary 29.
Let $\mathbbm{A}$ be a finite-dimensional algebra over some field $\mathbbm{F}$ and $n\in\mathbbm{N}$. Then every semiprime left ideal $I$ of $\mathbbm{A}[x_{1},\dots,x_{n}]$ is the intersection of all left ideals of $\mathbbm{A}[x_{1},\dots,x_{n}]$ of the form $J_{\theta,\xi,v}$ as in (16) that contain $I$.
Proof.
Let $I$ be a semiprime left ideal of $\mathbbm{A}[x_{1},\dots,x_{n}]$. By Theorem 27 and Proposition 28, $I$ is the intersection of some left ideals of the form $J_{\theta,\xi,v}$ as in (16), and therefore $I$ is the intersection of all left ideals of the form $J_{\theta,\xi,v}$ as in (16) that contain $I$. ∎